I was only able to work this problem out so far but I'm stuck.
I need to convert a quadratic equation to a vertex form.
y=-2x^2+2x+2
y-2=-2x^2+2x
y-2=-2(x^2-x)
If you could help me finish converting it and explain each step it'd be appreciated.

Answer by rothauserc(4718)

The equation of a parabola can be expressed in either standard or vertex form
---------------
standard form
y = ax^2 +bx +c
---------------
vertex form
y = a(x-h)^2 + k
---------------
we are given the standard form
y = -2x^2 +2x +2
the x coordinate for the vertex is
x = -b/2a = -2 / (2*(-2)) = (1/2)
the y coordinate of the vertex is
y = (-2*(1/2)^2) +(2*(1/2)) +2
y = -(1/2) + 1 +2 = 2.5
therefore the vertex is (.5, 2.5) and h = .5, k = 2.5
------------------
the vertex form is
y = -2(x-.5)^2 + 2.5